### v1

@@ 5 Conclusion
$\tau$ 0.275 [0.177 ; 0.357] 0.000 14,622 0.151
Panel (b): Heterogeneity with Discretize Income
@@
    Columns “RBC 95% CI” and “RBC $p$-value” report robust bias-corrected 95% confidence intervals and $p$-values, respectively, implemented using the data-driven bandwidth given in Column “$h$”. Variance Estimators are cluster-robust at the municipality level (as in the original study, Akhtari et al., 2022, column 1 of Table A.21). For more details, see Theorem 3, and the supplemental appendix.
@@
    For more details, see the supplemental appendix, and the replication files at https://rdpackages.github.io/replication/.
@@
    For more details, see the footnote of Table 1, and the replication files at https://rdpackages.github.io/replication/.

### v3

@@ 5 Conclusion
$\tau$ 0.275 [0.176 ; 0.357] 0.000 14,622 0.151
Panel (b): Heterogeneity with Discretized Income
@@
    Columns “RBC 95% CI” and “RBC $p$-value” report robust bias-corrected 95% confidence intervals and $p$-values, respectively, implemented using the data-driven bandwidth given in Column “$h$”. Variance Estimators are cluster-robust at the municipality level (as in the original study, Akhtari et al., 2022, column 1 of Table A.21). See Theorem 3, and the supplemental appendix, for more details.
@@
    See the replication files at https://rdpackages.github.io/replication/, and the supplemental appendix, for more details.
@@
    See the footnote of Table 1, and the replication files at https://rdpackages.github.io/replication/, for more details.